### v3

# Dimension-free uniform concentration bound for logistic regression

Abstract
We provide a novel dimension-free uniform concentration bound for the empirical risk function of constrained logistic regression. Our bound yields a milder sufficient condition for a uniform law of large numbers than conditions derived by the Rademacher complexity argument and McDiarmid’s inequality. The derivation is based on the PAC-Bayes approach with second-order expansion and Rademacher-complexity-based bounds for the residual term of the expansion.
Keywords: Effective rank, logistic regression, uniform concentration bound, uniform law of large numbers
MSC2020 Subject Classification: Primary 62J12; secondary 62F12
1 Introduction
The logistic regression model is a fundamental binary classification model in statistics and machine learning. Given a sequence of $\mathbb{R}^{p}\times\{0,1\}$-valued independent and identically distributed (i.i.d.) random variables $\{(\boldsymbol{X}_{i},Y_{i});i=1,\ldots,n\}$, it supposes $Y_{i}$ as conditionally Bernoulli-distributed random variables such that for some $\boldsymbol{\theta}\in\mathbb{R}^{p}$, for all $i=1,\ldots,n$,
$Y_{i}|\boldsymbol{X}_{i}=\boldsymbol{x}_{i}\sim\mathrm{Ber}(\sigma(\langle\boldsymbol{x}_{i},\boldsymbol{\theta}\rangle)),$ (1)
where $\sigma(t)=1/(1+\exp(t))$ with $t\in\mathbb{R}$ is the link function. Each component of $\boldsymbol{\theta}$ explains the relationship between the corresponding component of $\boldsymbol{X}_{i}$ and the conditional probability $\mathbb{P}(Y_{i}=1|\boldsymbol{X}_{i})$. The model is widely used for academic and industrial purposes as its interpretations are simple.
Our interest is the estimation of $\boldsymbol{\theta}$ with good prediction performance under high-dimensional settings. To estimate $\boldsymbol{\theta}$, we frequently consider the minimization problem of the following empirical risk function (or the $(-1/n)$-scaled log-likelihood function):
$\mathcal{R}_{n}(\boldsymbol{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\left(-Y_{i}\log\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{\theta}\rangle)-(1-Y_{i})\log(1-\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{\theta}\rangle)\right).$ (2)
We expect that the minimization of $\mathcal{R}_{n}(\boldsymbol{\theta})$ is a good approximation of the minimization of the unknown population risk function $\mathcal{R}(\boldsymbol{\theta}):=\mathbb{E}[\mathcal{R}_{n}(\boldsymbol{\theta})]$. If it holds true, the minimizers of $\mathcal{R}_{n}(\boldsymbol{\theta})$ also achieve small population risk and thus good prediction performance. Under the low-dimensional setting where $p$ is fixed in $n$ and $n\to\infty$, the classical argument for maximum likelihood estimation indeed validates this idea. However, Sur and Candès, (2019) point out that the minimizers of the empirical risk $\mathcal{R}_{n}(\boldsymbol{\theta})$ can perform poorly as the estimators of the minimizers of the population risk $\mathcal{R}(\boldsymbol{\theta})$ under high-dimensional settings. In contrast, our study examines when the minimization of $\mathcal{R}_{n}(\boldsymbol{\theta})$ is a good approximation of the minimization of $\mathcal{R}(\boldsymbol{\theta})$ even with large $p$.
In particular, we study uniform concentration bounds and a uniform law of large numbers as their corollary for $\mathcal{R}_{n}(\boldsymbol{\theta})$ around $\mathcal{R}(\boldsymbol{\theta})$ on $B[R]$, where $B[R]=\{\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{p};\|\boldsymbol{\theta}^{\prime}\|_{2}\leq R\}$ with $R\geq 0$ is the known bounded parameter space. Let us consider the following ball-constrained minimization problem (ball-constrained logistic regression):
minimize $\text{minimize }\mathcal{R}_{n}\left(\boldsymbol{\theta}\right)\text{ subject to }\left\|\boldsymbol{\theta}\right\|_{2}\leq R.$ (3)
This is a convex optimization problem on a bounded convex set; it has solutions, which we can find efficiently. If we can conclude that $\mathcal{R}_{n}(\boldsymbol{\theta})$ is uniformly close to $\mathcal{R}(\boldsymbol{\theta})$ on $B[R]$, then solving the minimization problem (3) is a good approximation of the minimization of $\mathcal{R}(\boldsymbol{\theta})$ on $B[R]$. The following uniform law of large numbers, one of the most fundamental arguments in large-sample theory, asymptotically supports this idea, provided that it holds:
$\lim_{n\to\infty}\sup_{\boldsymbol{\theta}\in B[R]}\left|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})\right|=0\text{ almost surely.}$. (4)
The uniform law (4) concludes that the minimization of the empirical risk $\mathcal{R}_{n}(\boldsymbol{\theta})$ is asymptotically equivalent to the minimization of the population risk $\mathcal{R}(\boldsymbol{\theta})$ (van der Vaart,, 2000). To derive sufficient conditions for the uniform law under high-dimensional setting, we analyse non-asymptotic uniform concentration bounds on $\sup_{\boldsymbol{\theta}\in B[R]}|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})|$.
A reasonable hypothesis is that $\mathcal{R}_{n}(\boldsymbol{\theta})$ is uniformly close to $\mathcal{R}(\boldsymbol{\theta})$ if an intrinsic dimension of the problem is sufficiently small in comparison to $n$. A promising intrinsic dimension is the effective rank of $\Sigma:=\mathbb{E}[\boldsymbol{X}_{1}\boldsymbol{X}_{1}^{\top}]$ defined as $\mathbf{r}(\Sigma):=\mathrm{tr}(\Sigma)/\|\Sigma\|$ as long as $\|\Sigma\|>0$, where $\|\Sigma\|$ is the spectral norm of $\Sigma$. The effective rank often appears in recent studies on possibly infinite-dimensional statistical models (Koltchinskii and Lounici,, 2017; Bartlett et al.,, 2020; Zhivotovskiy,, 2024). The classical Rademacher complexity argument and McDiarmid’s inequality (e.g., Bartlett and Mendelson,, 2002; Wainwright,, 2019; Bach,, 2023) indeed yield a dimension-free uniform concentration bound such that for all $\delta\in(0,1]$, with probability at least $1-\delta$,
$\sup_{\boldsymbol{\theta}\in B[R]}\left|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})\right|\leq 2\sqrt{\frac{R^{2}\|\Sigma\|\mathbf{r}(\Sigma)}{n}}+\sqrt{\frac{8\left(1+R^{2}K^{2}\|\Sigma\|\mathbf{r}(\Sigma)\right)\log\delta^{-1}}{n}}$ (5)
under a boundedness assumption $\|\boldsymbol{X}_{i}\|^{2}\leq K^{2}\mathrm{tr}(\Sigma)$ almost surely for some $K>0$. Note that $p$ does not appear on the right-hand side; therefore, this bound is dimension-free. We observe that the effective rank here explains an intrinsic dimension of the problem better than $p$ itself.
[…]
Our study gives a dimension-free uniform concentration bound yielding a milder sufficient condition for the uniform law of large numbers. We derive a bound such that for some explicit $c>0$ independent of $\Sigma$, $n$, and $p$, for any $\delta>0$, with probability at least $1-\delta$,
$\sup_{\boldsymbol{\theta}\in B[R]}\left|\mathcal{R}_{n}\left(\boldsymbol{\theta}\right)-\mathcal{R}\left(\boldsymbol{\theta}\right)\right|\leq c\sqrt{\frac{1+\|\Sigma\|\mathbf{r}(\Sigma)+(1+\|\Sigma\|)^{2}\log\delta^{-1}}{n}}.$ (6)
It is noteworthy that this bound gives a milder and more natural sufficient condition $\mathbf{r}(\Sigma)/n\to 0$ for the uniform law of large numbers than the classical bound (5).
The proof of our result is based on the PAC-Bayes approach with second-order expansion; e.g., see Catoni, (2007) and Alquier, (2024) for introductions to the PAC-Bayes approach. The PAC-Bayes approach is applicable to dimension-free analysis for linear problems such as covariance estimation (e.g., Zhivotovskiy,, 2024) and linear regression with least square losses, where functions of interests and their posterior means can coincide. We extend such approaches to nonlinear functions of interest (2) by considering the second-order expansion of risk functions. It generalizes previous studies for linear problems via the PAC-Bayes approach, as second-order terms are independent of parameters or even disappear under linearity.
1.1 Literature review
1.1.1 High-dimensional generalized linear models
Statistical inference for generalized linear models (GLMs) with large $p$ has been extensively studied. There are four typical settings on $\boldsymbol{\theta}$: moderately high-dimensional $\boldsymbol{\theta}$ with small $p/n$ (Liang and Du,, 2012; Hsu and Mazumdar,, 2023; Kuchelmeister and van de Geer,, 2024); sparse and high-dimensional $\boldsymbol{\theta}$ (van de Geer,, 2008; James and Radchenko,, 2009; Levy and Abramovich,, 2023); non-sparse and proportionally high-dimensional $\boldsymbol{\theta}$, i.e., $p/n\to\kappa\in(0,\infty)$ (Sur and Candès,, 2019; Sur et al.,, 2019; Salehi et al.,, 2019; Aubin et al.,, 2020; Emami et al.,, 2020; Sawaya et al.,, 2023); and non-sparse and possibly infinite-dimensional $\boldsymbol{\theta}$ (Wu et al.,, 2023). Note that our study adopts the last setting.
Liang and Du, (2012) discuss the asymptotic properties of the maximum likelihood estimation of logistic regression with $p=o(n)$. Hsu and Mazumdar, (2023) study the non-asymptotic sample complexities of constrained logistic regression with several inverse temperature settings and small $p/n$. Kuchelmeister and van de Geer, (2024) examine the non-asymptotic rates of convergence in the maximum likelihood estimation of constrained logistic regression under large and small signal-to-noise settings with small $p/n$. van de Geer, (2008) and James and Radchenko, (2009) study LASSO and Dantzig-selector estimators for GLMs with sparsity respectively. Levy and Abramovich, (2023) analyse the multinominal logistic regression problem under sparsity and show its generalization error. Sur and Candès, (2019) and Sur et al., (2019) investigate logistic regression under proportionally high-dimensional settings via the approximate message passing theory, show the biased behaviours of the ordinary maximum likelihood estimation, and propose procedures for debiased estimation and hypothesis testing. Arsov et al., (2019) study stability and generalization errors for some models including logistic regression. Salehi et al., (2019) investigate the impact of some regularization schemes on logistic regression under proportionally high-dimensional settings via the convex Gaussian min-max theorem. Aubin et al., (2020) consider some regularized empirical risk minimization problems including regularized logistic regression with proportionally high-dimensional settings and its Bayes optimality. Emami et al., (2020) analyse the asymptotic generalization errors of GLMs under proportional high-dimensional limits solved by a version of the approximate message passing algorithm. Sawaya et al., (2023) discuss the estimation of GLMs with general link functions with monotonicity and their asymmetric inverses. Wu et al., (2023) analyse GLMs with non-sparse and possibly infinite-dimensional $\boldsymbol{\theta}$ and the ReLU activation function as its link function trained by several algorithms.
1.1.2 Concentration of infinite-dimensional statistical models
Dimension-free bounds for non-sparse and possibly infinite-dimensional statistical models are a topical problem; note that most of the studies are based on linearity. Catoni and Giulini, (2017) consider the PAC-Bayes approach for estimation problems with possibly infinite-dimensional settings and heavy-tailed distributions and give several dimension-free bounds. Bartlett et al., (2020) show that the excess risk of the minimum $\ell^{2}$-norm interpolator for a linear regression model with $p>n$ gets small under the moderate decay of the eigenvalues of the covariance matrix of $\boldsymbol{X}_{i}$’s. Tsigler and Bartlett, (2023) further study ridge regression with $p>n$ in detail and give tighter bounds for the excess risk than Bartlett et al., (2020). Cheng and Montanari, (2022) also investigate ridge regression and develop a unified non-asymptotic analysis applicable to both underparameterized ($p<n$) and overparameterized ($p>n$) schemes.
1.2 Notation
For any vector $\boldsymbol{v}\in\mathbb{R}^{d}$, $\|\boldsymbol{v}\|$ denotes the $\ell^{2}$-norm of $\boldsymbol{v}$. For any vectors $\boldsymbol{u},\boldsymbol{v}\in\mathbb{R}^{d}$, $\langle\boldsymbol{u},\boldsymbol{v}\rangle$ represents the dot product of $\boldsymbol{u}$ and $\boldsymbol{v}$. For any matrix $A\in\mathbb{R}^{d_{1}}\otimes\mathbb{R}^{d_{2}}$, $\|A\|$ denotes the spectral norm of $A$ and $A^{\top}$ denotes the transpose of $A$. For any square matrix $A\in\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathrm{tr}(A)$ is the trace of $A$. For any positive semi-definite matrix $A\in\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathbf{r}(A)$ is the effective rank of $A$ defined as
$\mathbf{r}(A)=\begin{cases}\frac{\mathrm{tr}(A)}{\|A\|}&\text{ if }\|A\|>0,\\ 0&\text{ otherwise.}\end{cases}$ (7)
Let $B^{d}[r]:=\{\boldsymbol{\theta}\in\mathbb{R}^{d};\|\boldsymbol{\theta}\|\leq r\}$ with $r\geq 0$; we also use the notation $B[r]=B^{d}[r]$ if $d$ is obvious by the context. We define $\mathbb{S}^{d-1}:=\{\boldsymbol{u}\in\mathbb{R}^{d};\|\boldsymbol{u}\|=1\}$ with $d\in\mathbb{N}$.
For any function $f:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$, $f\in\mathcal{C}^{r}(\mathbb{R}^{d_{1}};\mathbb{R}^{d_{2}})$ means that $f$ is $r$-times continuously differentiable, $f\in\mathcal{C}_{b}^{r}(\mathbb{R}^{d_{1}};\mathbb{R}^{d_{2}})$ means that $f\in\mathcal{C}^{r}(\mathbb{R}^{d_{1}};\mathbb{R}^{d_{2}})$ and $f$ as well as its derivatives of order up to $r$ are bounded, and $f\in\mathcal{C}_{p}^{r}(\mathbb{R}^{d_{1}};\mathbb{R}^{d_{2}})$ means that $f\in\mathcal{C}^{r}(\mathbb{R}^{d_{1}};\mathbb{R}^{d_{2}})$ and $f$ as well as its derivatives of order up to $r$ are of at most polynomial growth. For any $f\in\mathcal{C}^{1}(\mathbb{R}^{d};\mathbb{R})$, $\nabla f:=(\partial_{1}f,\cdots,\partial_{d}f)$ is the gradient of $f$. For any $f\in\mathcal{C}^{2}(\mathbb{R}^{d};\mathbb{R})$, $\Delta f:=\sum_{i=1}^{d}\partial_{i}^{2}f$ denotes the Laplacian of $f$.
For functions $f,g:\mathbb{N}\to[0,\infty)$, $f(n)=\mathcal{O}(g(n))$ implies that for some $M>0$ and $n_{0}\in\mathbb{N}$, for all $n\geq n_{0}$, $f(n)/g(n)\leq M$, $f(n)=o(g(n))$ implies that $f(n)/g(n)\to 0$, $f(n)=\omega(g(n))$ implies that $f(n)/g(n)\to\infty$, and $f(n)\asymp g(n)$ implies that both $f(n)=\mathcal{O}(g(n))$ and $g(n)=\mathcal{O}(f(n))$ hold.
For any real-valued random variable $X$, $\|X\|_{\psi_{2}}:=\inf\{t>0;\mathbb{E}[\exp(X^{2}/t^{2})]\leq 2\}$, which is the $2$-Orlicz norm of $X$. For any $\mathbb{R}^{d}$-valued random variables $\boldsymbol{X}$, $\|\boldsymbol{X}\|_{\psi_{2}}:=\sup_{\boldsymbol{u}\in\mathbb{S}^{d-1}}\|\langle\boldsymbol{X},\boldsymbol{u}\rangle\|_{\psi_{2}}$ For two probability measures $\rho,\mu$ defined on an identical measurable space such that $\rho$ is absolutely continuous with respect to $\mu$, $H(\rho\|\mu)$ denotes the Kullback–Leibler divergence of $\rho$ from $\mu$. For two random variables $X,Y$ defined on an identical measurable space such that the law of $X$ is absolutely continuous with respect to that of $Y$, $H(X\|Y)$ denotes the Kullback–Leibler divergence of the law of $X$ from that of $Y$.
2 Main results
We begin with preparatory definition and assumption in Section 2.1 and give our main theorem, a dimension-free uniform concentration bound, in Section 2.2 and its corollary, a uniform law of large numbers, in Section 2.3. Section 2.4 presents an outline of the proof of the main theorem.
2.1 Assumption
We display the definition of the concentration property of random vectors and our assumption on $\{(\boldsymbol{X}_{i},Y_{i})\}$ before presenting the main results of this study.
Definition 1.
For any $d\in\mathbb{N}$, an $\mathbb{R}^{d}$-valued random variable $\boldsymbol{\zeta}$ satisfies the concentration property with constant $K>0$ if for every $1$-Lipschitz function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$, it holds that $\mathbb{E}[|\varphi(\boldsymbol{\zeta})|]<\infty$ and for all $u>0$,
$\mathbb{P}\left(\left|\varphi(\boldsymbol{\zeta})-\mathbb{E}[\varphi(\boldsymbol{\zeta})]\right|\geq u\right)\leq 2\exp\left(-u^{2}/K^{2}\right).$ (8)
The definition of the concentration property is adopted from Adamczak, (2015); Ledoux, (2001) gives a comprehensive overview of this type of concentration. A well-known sufficient condition for the concentration property is that the law of $\boldsymbol{Z}$ satisfies a logarithmic Sobolev inequality (see Section 5.4.2 of Bakry et al.,, 2014). For example, the $d$-dimensional standard Gaussian random vector satisfies the concentration property with constant $\sqrt{2}$ for arbitrary $d\in\mathbb{N}$ (see also Section 5.4 of Boucheron et al.,, 2013).
Assumption 1.
$\{(\boldsymbol{X}_{i},Y_{i});i=1,\ldots,n\}$ is a sequence of $\mathbb{R}^{p}\times\{0,1\}$-valued i.i.d. random variables with $\Sigma:=\mathbb{E}[\boldsymbol{X}_{i}\boldsymbol{X}_{i}^{\top}]$. For each $i=1,\ldots,n$, $\boldsymbol{X}_{i}$ has the representation $\boldsymbol{X}_{i}=U\Lambda^{1/2}\boldsymbol{Z}_{i}$, where $U$ is a $p\times p$ orthogonal matrix, $\Lambda$ is a $p\times p$ positive semi-definite diagonal matrix, and $\{\boldsymbol{Z}_{i}\}$ is a sequence of $p$-dimensional i.i.d. random vectors with the following properties:
  • (i)
    $\Sigma=U\Lambda U^{\top}$.
  • (ii)
    $\boldsymbol{Z}_{i}$ is isotropic, that is, $\mathbb{E}[\boldsymbol{Z}_{i}\boldsymbol{Z}_{i}^{\top}]=I_{p}$.
  • (iii)
    The $\mathbb{R}^{np}$-valued random variable $\boldsymbol{Z}:=(\boldsymbol{Z}_{1},\cdots,\boldsymbol{Z}_{n})$ satisfies the concentration property with constant $K>0$.
Note that we do not suppose $Y_{i}|\boldsymbol{X}_{i}=\boldsymbol{x}_{i}\sim\mathrm{Ber}(\sigma(\langle\boldsymbol{x}_{i},\boldsymbol{\theta}^{\ast}\rangle))$ for some $\boldsymbol{\theta}^{\ast}\in\mathbb{R}^{p}$ and thus we consider a possibly misspecified logistic regression problem.
Remark 1.
If Assumption 1 holds and $\mathbb{E}[\boldsymbol{Z}_{i}]=0$, then $\|\boldsymbol{Z}_{i}\|_{\psi_{2}}\leq\sqrt{3}K$, which can be checked as follows: for every $\boldsymbol{u}\in\mathbb{S}^{p-1}$,
$\mathbb{E}\left[\exp\left(\langle\boldsymbol{Z}_{i},\boldsymbol{u}\rangle^{2}/(3K^{2})\right)\right]=1+\int_{1}^{\infty}\mathbb{P}\left(|\langle\boldsymbol{Z}_{i},\boldsymbol{u}\rangle|\geq\sqrt{3K^{2}\log t}\right)\mathrm{d}t\leq 1+2\int_{1}^{\infty}t^{-3}\mathrm{d}t=2.$
Otherwise, $\|\boldsymbol{Z}_{i}\|_{\psi_{2}}\leq\|\boldsymbol{Z}_{i}-\mathbb{E}[\boldsymbol{Z}_{i}]\|_{\psi_{2}}+\|\mathbb{E}[\boldsymbol{Z}_{i}]\|_{\psi_{2}}\leq\sqrt{3}K+\sup_{\boldsymbol{u}\in\mathbb{S}^{p-1}}|\mathbb{E}[\langle\boldsymbol{Z}_{i},\boldsymbol{u}\rangle]|\leq\sqrt{3}K+\|\mathbb{E}[\boldsymbol{Z}_{i}\boldsymbol{Z}_{i}^{\top}]\|^{1/2}\leq 1+\sqrt{3}K$.
2.2 Main theorem
We show a dimension-free uniform concentration inequality for the empirical risk function of logistic regression, which is the main theorem of this study.
Theorem 1.
Suppose that Assumption 1 holds. For arbitrary $\delta\in(0,1/6]$, $n\in\mathbb{N}$, and $R\geq 0$ such that $n>\|\Sigma\|(1+6R^{2})(1+\sqrt{3}K)^{2}/12$, with probability at least $1-6\delta$,
$\sup_{\boldsymbol{\theta}\in B[R]}\left|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})\right|$ $\leq\sqrt{\frac{27(\log\delta^{-1}+(1+\sqrt{3}K)^{2}(\mathrm{tr}(\Sigma)/12+\|\Sigma\|R^{2}\log\delta^{-1}))(1+6R^{2})}{n}}$
$\quad+2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}}+\sqrt{\frac{78K^{2}\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)}{n}}+\frac{9K^{2}R\|\Sigma\|\sqrt{\log\delta^{-1}}}{n}.$ (9)
The proof is given in Section 4.2; we give a sketch of the proof in Section 2.4.
The characteristics of our bounds are its dimension-free property and explicit representation of numerical constants. The bound given by Theorem 1 is only dependent on $\delta,K,\mathrm{tr}(\Sigma),\|\Sigma\|$, and $R$, and thus dimension-free. Hence, we can extend our result for $\mathbb{R}^{p}$ to general separable Hilbert spaces by the dimension-free property and the monotonicity of probability measures (for example, see Giulini,, 2018). Note that every numerical constant is explicit; it is sometimes not easy to obtain such explicit constants in dimension-free analysis (e.g., Koltchinskii and Lounici,, 2017). Explicit representation enables us to give uncertainty quantification of the population risk at a solution of the constrained minimization problem (3) provided that $K$ is known and deterministic or high-probability upper bounds for $\|\Sigma\|$ and $\mathrm{tr}(\Sigma)$ are available.
2.3 Uniform law of large numbers
We consider a sufficient condition for the uniform law of large numbers (4) via Theorem 1. To deal with $p$ and $\Sigma$ possibly dependent on $n$, we define a double array of random variables satisfying Assumption 1 for each $n$.
Assumption 2.
For each $n\in\mathbb{N}$, $p_{n}\in\mathbb{N}$ and $\{(\boldsymbol{X}_{i,n},Y_{i,n});i=1,\ldots,n\}$ is a sequence of $\mathbb{R}^{p_{n}}\times\{0,1\}$-valued i.i.d. random variables with $\Sigma_{n}:=\mathbb{E}[\boldsymbol{X}_{1,n}\boldsymbol{X}_{1,n}^{\top}]$. For some $K>0$, $\boldsymbol{X}_{i,n}=U_{n}\Lambda_{n}^{1/2}\boldsymbol{Z}_{i,n}$ holds for all $i=1,\ldots,n$ and $n\in\mathbb{N}$, where $U_{n}$ is a $p_{n}\times p_{n}$ orthogonal matrix, $\Lambda_{n}$ is a $p_{n}\times p_{n}$ positive semi-definite diagonal matrix, and $\{\boldsymbol{Z}_{i,n};i=1,\ldots,n\}$ is a sequence of $p_{n}$-dimensional i.i.d. random vectors with the following properties: (i) $\Sigma_{n}=U_{n}\Lambda_{n}U_{n}^{\top}$; (ii) $\mathbb{E}[\boldsymbol{Z}_{i,n}\boldsymbol{Z}_{i,n}^{\top}]=I_{p_{n}}$; (iii) $\boldsymbol{Z}_{n}:=(\boldsymbol{Z}_{1,n},\cdots,\boldsymbol{Z}_{n,n})$ satisfies the concentration property with constant $K$.
We define the empirical risk function under Assumption 2 using $\{(\boldsymbol{X}_{i,n},Y_{i,n})\}$ instead of $\{(\boldsymbol{X}_{i},Y_{i})\}$ for each $n$ as follows: for each $\boldsymbol{\theta}\in\mathbb{R}^{p_{n}}$,
$\mathcal{R}_{n}(\boldsymbol{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\left(-Y_{i,n}\log\sigma(\langle\boldsymbol{X}_{i,n},\boldsymbol{\theta}\rangle)-(1-Y_{i,n})\log(1-\sigma(\langle\boldsymbol{X}_{i,n},\boldsymbol{\theta}\rangle)\right).$ (10)
The following corollary yields a sufficient condition of the uniform law of large numbers; it holds by Theorem 1 with letting $\delta=1/n^{2}$ and the Borel–Cantelli Lemma.
Corollary 2 (Uniform law of large numbers).
Suppose that Assumption 2 holds and $\sup_{n\in\mathbb{N}}\|\Sigma_{n}\|<\infty$. Fix an arbitrary $R\geq 0$. If $\lim_{n\to\infty}\mathbf{r}(\Sigma_{n})/n=0$, then
$\lim_{n\to\infty}\sup_{\boldsymbol{\theta}\in B^{p_{n}}[R]}\left|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathbb{E}[\mathcal{R}_{n}(\boldsymbol{\theta})]\right|=0\text{ almost surely.}$. (11)
We can also know that the minimum empirical risk converges to the minimum population risk almost surely if $\mathbf{r}(\Sigma_{n})/n\to 0$ by Corollary 2. This is milder than reasonable sufficient conditions derived by classical arguments; we discuss it in Section 3.
2.4 Idea of the proof
We exhibit the idea of the proof of Theorem 1 briefly. We first use the PAC-Bayes approach by adopting $\{\boldsymbol{W}_{t}^{\boldsymbol{\theta}};t\geq 0\}$ with $\boldsymbol{\theta}\in B[R]$ , $p$-dimensional Brownian motions starting at $\boldsymbol{\theta}$, as the prior and posteriors and see that for any $t\in(0,1]$, with high probability, for all $\boldsymbol{\theta}\in B[R]$,
$\mathbb{E}_{\boldsymbol{W}}\left[\mathcal{R}\left(\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]-\mathbb{E}_{\boldsymbol{W}}\left[\mathcal{R}_{n}\left(\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]\leq\text{(PAC-Bayes complexity bound for $\mathcal{R}_{n}$)}.$ (12)
In the second place, we use Itô’s formula and obtain that for any $t\in(0,1]$, with high probability, for all $\boldsymbol{\theta}\in B[R]$,
[…]
$\quad+\frac{1}{2}\int_{0}^{t}\left(\mathbb{E}_{\boldsymbol{W}}\left[\Delta\mathcal{R}_{n}\left(\boldsymbol{W}_{s}^{\boldsymbol{\theta}}\right)\right]-\mathbb{E}_{\boldsymbol{W}}\left[\Delta\mathcal{R}\left(\boldsymbol{W}_{s}^{\boldsymbol{\theta}}\right)\right]\right)\mathrm{d}s.$ (13)
Note that the Laplacian of $\mathcal{R}_{n}$ is no more dependent on $\{Y_{i}\}$ and we can reduce the problem of $\{(\boldsymbol{X}_{i},Y_{i})\}$ to that of $\{\boldsymbol{X}_{i}\}$ or equivalently $\{\boldsymbol{Z}_{i}\}$, whose behaviours are tractable by Assumption 1. We thirdly give a Bernstein-type inequality of the supremum of the second term on the right-hand side in $\boldsymbol{\theta}$ around its expectation in $\boldsymbol{X}$ and derive a bound such that for any $t\in(0,1]$, with high probability, for all $\boldsymbol{\theta}\in B[R]$,
[…]
In the last place, we evaluate a Rademacher-complexity-based bound for the second term on the right-hand side. Since the sign on the left-hand side does not affect the discussion, we complete the proof by selecting appropriate $t\in(0,1]$.
Remark 2.
When we consider linear problems such as covariance estimation (e.g., Zhivotovskiy,, 2024) and linear regression with least square losses, Laplacians are independent of parameters or even equal to zero. We thus can regard our approach as a natural extension of the PAC-Bayes approach for linear problems to nonlinear ones.
Remark 3.
We can yield similar conclusions via Taylor’s theorem instead of Itô’s formula to expand the expectation with respect to Gaussian vectors; however, Itô’s formula gives simpler proofs than Taylor’s theorem. One of the reasons is that we repeatedly employ Stein’s lemma and the Hermite polynomials in our proofs. Itô’s formula leads to the remainder with the integrand of the tractable form $\sigma(1-\sigma)H_{0}$, where $H_{d}$ with $d\in\mathbb{N}_{0}$ is the $d$-th Hermite polynomial. On the other hand, Taylor’s theorem gives the remainder with the integrand of the form $\sigma(1-\sigma)(H_{2}+H_{0})$ resulting in lengthy proofs.
3 Discussions
We compare our result with the classical argument by Rademacher complexities and McDiarmid’s inequality and its extension to unbounded cases from the viewpoint of the uniform law of large numbers. Dealing with Gaussian $\boldsymbol{X}_{i}$ as previous studies, one may consider an extension of the classical argument by the concentration of anisotropic subgaussian random vectors other than our approach; hence, we here give such an extension for comparison. In this section, we suppose that Assumption 2 holds. Furthermore, assume that each $\boldsymbol{Z}_{i,n}$ is centred and its components are independent.
A uniform concentration bound (5) is of order $\mathcal{O}(\sqrt{\mathbf{r}(\Sigma_{n})\log\delta^{-1}/n})$, and an extension of the classical argument to subgaussian cases also gives the same order. Theorem 6.3.2 of Vershynin, (2018) gives that there exists an absolute constant $a>0$ such that for any $t\geq 0$,
$\mathbb{P}\left(\max_{i=1,\ldots,n}\left|\|U_{n}\Lambda_{n}^{1/2}\boldsymbol{Z}_{i,n}\|-\mathrm{tr}(\Sigma_{n})^{1/2}\right|\geq a\sqrt{K^{4}\|\Sigma_{n}\|^{2}(t+\log n)}\right)\leq 2\exp(-t).$ (15)
Since $|-y\log\sigma(\langle x,\boldsymbol{\theta}\rangle)-(1-y)\log(1-\sigma(\langle x,\boldsymbol{\theta}\rangle))|^{2}\leq 2(1+R^{2}\|x\|^{2})$ for all $y\in\{0,1\}$, $x\in\mathbb{R}^{p_{n}}$, and $\boldsymbol{\theta}\in B^{p_{n}}[R]$, with probability at least $1-2\delta$, it holds that $\max_{i=1,\ldots,n}\sup_{\boldsymbol{\theta}\in\Theta}|-Y_{i}\log\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{\theta}\rangle)-(1-Y_{i})\log(1-\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{\theta}\rangle))|^{2}\leq 2(1+2R^{2}(\mathrm{tr}(\Sigma_{n})+a^{2}K^{4}\|\Sigma_{n}\|^{2}(\log n+\log\delta^{-1})))$. Therefore, the classical argument (see Section 4.5 of Bach,, 2023) gives that for arbitrary $\delta>0$, with probability $1-3\delta$,
$\sup_{\boldsymbol{\theta}\in B^{p_{n}}[R]}\left|\mathcal{R}_{n}(\boldsymbol{\theta})-\mathbb{E}[\mathcal{R}_{n}(\boldsymbol{\theta})]\right|$
$\leq 2\sqrt{\frac{R^{2}\|\Sigma_{n}\|\mathbf{r}(\Sigma_{n})}{n}}+\sqrt{\frac{8(1+2R^{2}\|\Sigma_{n}\|(\mathbf{r}(\Sigma_{n})+a^{2}K^{4}\|\Sigma_{n}\|(\log n+\log\delta^{-1})))\log\delta^{-1}}{n}}.$ (16)
The right-hand side above is of order $\mathcal{O}(\sqrt{\mathbf{r}(\Sigma_{n})\log\delta^{-1}/n})$, which is the same as the estimate (5).
Corollary 2 shows that a sufficient condition for the uniform law of large numbers (4) is $\mathbf{r}(\Sigma_{n})/n\to 0$, which is milder than ones based on the estimates (5) and (3). Both the estimates (5) and (3) combined with the Borel–Cantelli lemma and $\delta=1/n^{2}$ give a sufficient condition of the uniform law of large numbers $\mathbf{r}(\Sigma_{n})\log n/n\to 0$, which is worse than $\mathbf{r}(\Sigma_{n})/n\to 0$.
4 Proof of the main theorem
We first introduce three preliminary results in Section 4.1 and show the proof of Theorem 1 using them in Section 4.2. The proofs of the preliminary results are given in Appendix.
4.1 Preliminary results
4.1.1 PAC-Bayes bounds with second-order expansion
Let $(\mathbb{M},\mathcal{M})$ be a measurable space and $\{V_{i};i=1\ldots,n\}$ be a sequence of $(\mathbb{M},\mathcal{M})$-valued i.i.d. random variables with probability measure $P^{V}$. Fix $R\geq 0$. We let $\{\boldsymbol{W}_{t}^{\boldsymbol{\theta}};t\geq 0\}$ with $\boldsymbol{\theta}\in B^{p}[R]$ represent a $p$-dimensional Brownian motion starting at $\boldsymbol{\theta}$, that is, $\boldsymbol{W}_{t}^{\boldsymbol{\theta}}=\boldsymbol{\theta}+\boldsymbol{W}_{t}$, where $\{\boldsymbol{W}_{t};t\geq 0\}$ is the $p$-dimensional standard Brownian motion starting at zero.
Proposition 3 (PAC-Bayes bounds with second-order expansion).
Fix $t>0$. Let $f:\mathbb{M}\times\mathbb{R}^{p}\ni(v,\boldsymbol{w})\mapsto f(v,\boldsymbol{w})\in\mathbb{R}$ be a measurable function satisfying the following:
  • (i)
    For $P^{V}$-almost all $v\in\mathbb{M}$, the following hold: $f(v,\cdot)\in\mathcal{C}^{2}(\mathbb{R}^{p};\mathbb{R})$, and the gradient and Laplacian of $f(v,\cdot)$ equal some measurable functions $\nabla_{\boldsymbol{w}}f:\mathbb{M}\times\mathbb{R}^{p}\to\mathbb{R}^{p}$ and $\Delta_{\boldsymbol{w}}f:\mathbb{M}\times\mathbb{R}^{p}\to\mathbb{R}$ such that
    • (a)
      $\nabla_{\boldsymbol{w}}f(v,\cdot)$ is at most of polynomial growth.
    • (b)
      For all $\boldsymbol{\theta}\in B[R]$,
      $\int_{0}^{t}\mathbb{E}_{V,\boldsymbol{W}}\left[\left|\left.\Delta_{\boldsymbol{w}}f(V_{i},\boldsymbol{w})\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right|\right]\mathrm{d}s<\infty.$
  • (ii)
    For all $\boldsymbol{w}\in\mathbb{R}^{p}$, $|f(V_{i},\boldsymbol{w})|<\infty$ almost surely, and for all $\boldsymbol{\theta}\in B[R]$, $\mathbb{E}_{V}[|f(V_{i},\boldsymbol{\theta})|]<\infty$.
  • (iii)
    For some $\eta:\mathbb{R}^{p}\to[0,\infty)$, for all $\lambda\in\mathbb{R}$ and $\boldsymbol{w}\in\mathbb{R}^{p}$,
    $\mathbb{E}_{V}\left[\exp\left(\lambda\left(f(V_{1},\boldsymbol{w})-\mathbb{E}_{V}\left[f(V_{1},\boldsymbol{w})\right]\right)\right)\right]\leq\exp\left(\frac{\lambda^{2}\eta^{2}(\boldsymbol{w})}{8}\right),$
    and $\sup_{\boldsymbol{\theta}\in B[R]}\mathbb{E}[\eta^{2}(\boldsymbol{W}_{t}^{\boldsymbol{\theta}})]<\infty$.
Then, if $n\in\mathbb{N}$, $\delta\in(0,1]$, and $C_{t}\in[\sup_{\boldsymbol{\theta}\in B[R]}\mathbb{E}[\eta^{2}(\boldsymbol{W}_{t}^{\boldsymbol{\theta}})],\infty)$ satisfy
$\mathbb{E}_{V,\boldsymbol{W}}\left[\exp\left(\left(f\left(V_{1},\boldsymbol{W}_{t}\right)-\mathbb{E}_{V}\left[f\left(V_{1},\boldsymbol{W}_{t}\right)\right]\right)\sqrt{\frac{8(R^{2}/(2t)+\log\delta^{-1})}{nC_{t}}}\right)\right]<\infty,$ (17)
then with probability at least $1-\delta$, for all $\boldsymbol{\theta}\in B[R]$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{V}\left[f\left(V_{i},\boldsymbol{\theta}\right)\right]\leq$ $\,\frac{1}{n}\sum_{i=1}^{n}f\left(V_{i},\boldsymbol{\theta}\right)+\sqrt{\frac{C_{t}(R^{2}/(2t)+\log\delta^{-1})}{2n}}$
$+\frac{1}{2n}\sum_{i=1}^{n}\int_{0}^{t}\left(\mathbb{E}_{\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]-\mathbb{E}_{V,\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]\right)\mathrm{d}s,$ (18)
and with probability at least $1-\delta$, for all $\boldsymbol{\theta}\in B[R]$,
$\frac{1}{n}\sum_{i=1}^{n}f\left(V_{i},\boldsymbol{\theta}\right)\leq$ $\,\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{V}\left[f\left(V_{i},\boldsymbol{\theta}\right)\right]+\sqrt{\frac{C_{t}(R^{2}/(2t)+\log\delta^{-1})}{2n}}$
$+\frac{1}{2n}\sum_{i=1}^{n}\int_{0}^{t}\left(\mathbb{E}_{V,\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]-\mathbb{E}_{\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]\right)\mathrm{d}s.$ (19)
4.1.2 Bernstein-type inequalities
Let $\boldsymbol{z}=(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})\in\mathbb{R}^{np}$ and $\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}}=\Lambda^{1/2}U^{\top}(\boldsymbol{\theta}+\boldsymbol{W}_{s})$. We set
$G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z})$ $:=\frac{1}{2n}\sum_{i=1}^{n}\int_{0}^{t}\mathbb{E}_{\boldsymbol{W}}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\right]\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\mathrm{d}s$
$\quad-\frac{1}{2n}\sum_{i=1}^{n}\int_{0}^{t}\mathbb{E}_{\boldsymbol{W},\boldsymbol{Z}}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{Z}_{i}\right\rangle\right)\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle\right]\mathrm{d}s.$ (20)
We give the following Bernstein-type inequalities.
Proposition 4.
Suppose that Assumption 1 holds. For all $u>0$ and $t>0$,
$\mathbb{P}$ $\left(\left|\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})-\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})\right]\right|\geq u\right)$
$\leq 2\exp\left(-\frac{n}{6K^{2}}\min\left\{\frac{u^{2}}{155t\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)},\frac{u}{5\sqrt{t}R\|\Sigma\|}\right\}\right),$ (21)
and
$\mathbb{P}$ $\left(\left|\sup_{\boldsymbol{\theta}\in B[R]}(-1)G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})-\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}(-1)G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})\right]\right|\geq u\right)$
$\leq 2\exp\left(-\frac{n}{6K^{2}}\min\left\{\frac{u^{2}}{155t\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)},\frac{u}{5\sqrt{t}R\|\Sigma\|}\right\}\right).$ (22)
4.1.3 Expectation of the suprema of the Laplacian gaps
Proposition 5.
Suppose that Assumption 1 holds. For any $t\in(0,1]$,
$\mathbb{E}_{\boldsymbol{Z}}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}\left(\boldsymbol{Z}\right)\right]\leq 2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}},\quad\mathbb{E}_{\boldsymbol{Z}}\left[\sup_{\boldsymbol{\theta}\in B[R]}(-1)G_{t}^{\boldsymbol{\theta}}\left(\boldsymbol{Z}\right)\right]\leq 2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}}.$ (23)
4.2 Proof of Theorem 1
Proof of Theorem 1.
We discuss a uniform upper bound for $\mathcal{R}(\boldsymbol{\theta})-\mathcal{R}_{n}(\boldsymbol{\theta})$ and see that the same argument holds for $\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})$ immediately by flipping the sign.
(Step 1) Since $-y\log(\sigma(t))-(1-y)\log((1-\sigma(t)))\in[0,\log 2+|t|]$ for all $y\in\{0,1\}$ and $t\in\mathbb{R}$,
$f(V_{i},\boldsymbol{w}):=-Y_{i}\log(\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{w}\rangle))-(1-Y_{i})\log(1-\sigma(\langle\boldsymbol{X}_{i},\boldsymbol{w}\rangle))$ (24)
with $V_{i}=(\boldsymbol{X}_{i},Y_{i})$ (or equivalently $V_{i}=(\boldsymbol{Z}_{i},Y_{i})$) for each $\boldsymbol{w}\in\mathbb{R}^{p}$ is a subgaussian random variable such that $\|f(V_{i},\boldsymbol{w})\|_{\psi_{2}}\leq\log 2+\|\boldsymbol{Z}_{i}\|_{\psi_{2}}\|\Lambda^{1/2}U^{\top}\boldsymbol{w}\|$. Proposition 2.5.2 of Vershynin, (2018) (the flow (iv) ×$\Rightarrow^{\times 1}$ (i) ×$\Rightarrow^{\times 2}$ (ii), where $x>0$ in ×$\Rightarrow^{\times x}$ denotes the absolute scales in the proof) gives $\|f(V_{i},\boldsymbol{w})-\mathbb{E}_{V}[f(V_{i},\boldsymbol{w})]\|_{\psi_{2}}\leq 3\|f(V_{i},\boldsymbol{w})\|_{\psi_{2}}$. Using $\|\boldsymbol{Z}_{i}\|_{\psi_{2}}\leq 1+\sqrt{3}K$ by Assumption 1, Proposition 2.5.2 of Vershynin, (2018) (the flow (iv) ×$\Rightarrow^{\times 1}$ (i) ×$\Rightarrow^{\times 2}$ (ii) ×$\Rightarrow^{\times 1/(2\sqrt{e})}$ (iii) ×$\Rightarrow^{\times 1}$ (v)) gives that for any $\lambda\in\mathbb{R}$ and $\boldsymbol{w}\in\mathbb{R}^{p}$,
$\mathbb{E}_{V}\left[\exp\left(\lambda\left(f(V_{i},\boldsymbol{w})-\mathbb{E}_{V}[f(V_{i},\boldsymbol{w})]\right)\right)\right]\leq\exp\left(\frac{9\lambda^{2}}{e}\left(\log 2+\left(1+\sqrt{3}K\right)\left\|\Lambda^{1/2}U^{\top}\boldsymbol{w}\right\|\right)^{2}\right).$ (25)
Thus we choose $\eta:\mathbb{R}^{p}\to[0,\infty)$ such that
$\eta^{2}(\boldsymbol{w}):=\frac{72}{e}\left(\log 2+\left(1+\sqrt{3}K\right)\left\|\Lambda^{1/2}U^{\top}\boldsymbol{w}\right\|\right)^{2}.$ (26)
Since
$\mathbb{E}_{\boldsymbol{W}}\left[\left\|\Lambda^{1/2}U^{\top}\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right\|^{2}\right]=\mathbb{E}_{\boldsymbol{W}}\left[\left\langle\Sigma\left(\boldsymbol{W}_{t}+\boldsymbol{\theta}\right),\boldsymbol{W}_{t}+\boldsymbol{\theta}\right\rangle\right]=t\mathrm{tr}(\Sigma)+\langle\Sigma\boldsymbol{\theta},\boldsymbol{\theta}\rangle,$
for all $t>0$ and $\boldsymbol{\theta}\in B[R]$,
$\mathbb{E}_{\boldsymbol{W}}\left[\eta^{2}\left(\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]\leq\frac{144}{e}\left(1+\left(1+\sqrt{3}K\right)^{2}\left(t\mathrm{tr}(\Sigma)+\|\Sigma\|R^{2}\right)\right)=:C_{t}$ (27)
by the fact that $(a+b)^{2}\leq 2(a^{2}+b^{2})$ for all $a,b\in\mathbb{R}$.
(Step 2) We apply Proposition 3 for $\mathcal{R}(\boldsymbol{\theta})-\mathcal{R}_{n}(\boldsymbol{\theta})$. For all $t>0$, Conditions (i) and (ii) of Proposition 3 are immediate since the gradient and Laplacian of $f$ are
$\nabla_{\boldsymbol{w}}f(V_{i},\boldsymbol{w})$ $=-Y_{i}\boldsymbol{X}_{i}+\sigma\left(\left\langle\boldsymbol{X}_{i},\boldsymbol{w}\right\rangle\right)\boldsymbol{X}_{i}=-Y_{i}U\Lambda^{1/2}\boldsymbol{Z}_{i}+\sigma\left(\left\langle U\Lambda^{1/2}\boldsymbol{Z}_{i},\boldsymbol{w}\right\rangle\right)U\Lambda^{1/2}\boldsymbol{Z}_{i},$
$\Delta_{\boldsymbol{w}}f(V_{i},\boldsymbol{w})$ $=\left(\sigma\left(1-\sigma\right)\right)\left(\langle\boldsymbol{X}_{i},\boldsymbol{w}\rangle\right)\left\|\boldsymbol{X}_{i}\right\|^{2}=\left(\sigma\left(1-\sigma\right)\right)\left(\left\langle U\Lambda^{1/2}\boldsymbol{Z}_{i},\boldsymbol{w}\right\rangle\right)\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle.$
[…]
$\mathcal{R}(\boldsymbol{\theta})-\mathcal{R}_{n}(\boldsymbol{\theta})$ $\leq\sqrt{\frac{27(1+(1+\sqrt{3}K)^{2}(t\mathrm{tr}(\Sigma)+\|\Sigma\|R^{2}))(R^{2}/(2t)+\log(\delta^{-1}))}{n}}$
$\quad+\frac{1}{2n}\sum_{i=1}^{n}\int_{0}^{t}\left(\mathbb{E}_{\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]-\mathbb{E}_{V,\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{s}^{\boldsymbol{\theta}}}\right]\right)\mathrm{d}s.$ (28)
(Step 3) By Propositions 4 and 5, for all $t\in\mathcal{T}\subset(0,1]$, with probability at least $1-3\delta$, for all $\boldsymbol{\theta}\in B[R]$,
$\mathcal{R}(\boldsymbol{\theta})-\mathcal{R}_{n}(\boldsymbol{\theta})$ $\leq\sqrt{\frac{27(1+(1+\sqrt{3}K)^{2}(t\mathrm{tr}(\Sigma)+\|\Sigma\|R^{2}))(R^{2}/(2t)+\log(\delta^{-1}))}{n}}$
$\quad+2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}}+\sqrt{\frac{930tK^{2}\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)\log\delta^{-1}}{n}}+\frac{30\sqrt{t}K^{2}R\|\Sigma\|\log\delta^{-1}}{n}.$ (29)
If $t^{\ast}:=1/(12\log\delta^{-1})\in\mathcal{T}$ (note $\log\delta^{-1}>1$ by $\delta\leq 1/6$), by fixing $t=t^{\ast}$, we obtain that with probability at least $1-3\delta$, for all $\boldsymbol{\theta}\in B[R]$,
$\mathcal{R}(\boldsymbol{\theta})-\mathcal{R}_{n}(\boldsymbol{\theta})$ $\leq\sqrt{\frac{27(\log\delta^{-1}+(1+\sqrt{3}K)^{2}(\mathrm{tr}(\Sigma)/12+\|\Sigma\|R^{2}\log\delta^{-1}))(1+6R^{2})}{n}}$
$\quad+2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}}+\sqrt{\frac{78K^{2}\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)}{n}}+\frac{9K^{2}R\|\Sigma\|\sqrt{\log\delta^{-1}}}{n}.$ (30)
(Step 4) We now check if $t^{\ast}\in\mathcal{T}$. By noting that $C_{t}\geq 144/e$ for all $t>0$, we see that
$\lambda^{\ast}:=\sqrt{\frac{8(R^{2}/(2t^{\ast})+\log\delta^{-1})}{nC_{t^{\ast}}}}\leq\sqrt{\frac{e(1+6R^{2})\log\delta^{-1}}{18n}}.$ (31)
By Eq. (25),
$\mathbb{E}_{V,\boldsymbol{W}}\left[\exp\left(\lambda^{\ast}\left(f(V_{1},\boldsymbol{W}_{t^{\ast}})-\mathbb{E}_{V}\left[f(V_{1},\boldsymbol{W}_{t^{\ast}})\right]\right)\right)\right]$
$\leq\mathbb{E}_{\boldsymbol{W}}\left[\exp\left(\frac{(1+6R^{2})\log\delta^{-1}}{2n}\left(\log 2+\left(1+\sqrt{3}K\right)\left\|\Lambda^{1/2}U^{\top}\boldsymbol{W}_{t^{\ast}}\right\|\right)^{2}\right)\right].$ (32)
It suffices to show that the right-hand side is finite. For all $\alpha\in\mathbb{R}$, $\mathbb{E}_{\boldsymbol{W}}[\exp(\alpha\|\Lambda^{1/2}U^{\top}\boldsymbol{W}_{t}\|)]<\infty$. Moreover, for all $t>0$ and $\beta\in[0,1/(2\|\Sigma\|t))$, $\mathbb{E}_{\boldsymbol{W}}[\exp(\beta\|\Lambda^{1/2}U^{\top}\boldsymbol{W}_{t}\|^{2})]<\infty$. Hence, the right-hand side of (4.2) is finite if
$\frac{\left(1+6R^{2}\right)\left(1+\sqrt{3}K\right)^{2}\log\delta^{-1}}{2n}<\frac{6\log\delta^{-1}}{\|\Sigma\|}$ (33)
or equivalently $\|\Sigma\|(1+6R^{2})(1+\sqrt{3}K)^{2}/12<n$; this is assumed. Therefore, $t^{\ast}\in\mathcal{T}$ indeed.
(Step 5) The same argument holds for $\mathcal{R}_{n}(\boldsymbol{\theta})-\mathcal{R}(\boldsymbol{\theta})$ by applying Propositions 35.
Hence, we obtain the conclusion. ∎
Remark 4.
We select $t=1/(12\log\delta^{-1})$, which approximately solves the following quadratic equation on $(0,\infty)$ under $K=\sqrt{2}$ (for standard Gaussian distributions), $R\gg 1$, $R=\mathcal{O}(1)$, $\|\Sigma\|=\mathcal{O}(1)$, and $\mathrm{tr}(\Sigma)=\omega(1)$:
$1860t^{2}\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)\log\delta^{-1}=27(1+(1+\sqrt{6})^{2}(t\mathrm{tr}(\Sigma)+\|\Sigma\|R^{2}))(R^{2}/2+t\log(\delta^{-1})).$
This equation lets the first and the third terms on the right-hand side of (4.2) be equal to each other. Under different asymptotic settings, the optimal selection of $t$ can be different to our selection.
Acknowledgements
I gratefully acknowledge Pierre Alquier for his enlightening comments. This work was supported by JSPS KAKENHI Grant Number JP24K02904 and JST CREST Grant Numbers JPMJCR21D2 and JPMJCR2115.
Appendix A On the Hermite polynomials
We present some known results on Stein’s lemma and the Hermite polynomials used in the proofs of the technical results for readers’ convenience. Let us introduce the Hermite polynomials of degrees only up to 3: for any $x\in\mathbb{R}$, $H_{0}(x)=1$, $H_{1}(x)=x$, $H_{2}(x)=x^{2}-1$, and $H_{3}(x)=x^{3}-3x$.
Lemma 6.
Suppose that $\zeta\sim\mathcal{N}(0,1)$. The following hold:
  1. (i)
    For any $f\in\mathcal{C}_{p}^{1}(\mathbb{R};\mathbb{R})$ and $d=0,1,2$, $\mathbb{E}[f^{\prime}(\zeta)H_{d}(\zeta)]=\mathbb{E}[f(\zeta)H_{d+1}(\zeta)]$.
  2. (ii)
    For any $f\in\mathcal{C}_{p}^{2}(\mathbb{R};\mathbb{R})$ and $d=0,1$, $\mathbb{E}[f^{\prime\prime}(\zeta)H_{d}(\zeta)]=\mathbb{E}[f(\zeta)H_{d+2}(\zeta)]$.
Proof.
The statement (i) with $d=0$ is nothing more than Stein’s lemma. Since any $f,g\in\mathcal{C}_{p}^{1}(\mathbb{R};\mathbb{R})$,
$\mathbb{E}[f^{\prime}(\zeta)g(\zeta)]=\mathbb{E}[(fg)^{\prime}(\zeta)-f(\zeta)g^{\prime}(\zeta)]=\mathbb{E}[f(\zeta)(g(\zeta)\zeta-g^{\prime}(\zeta))],$
we can easily check that the other statements hold. ∎
Appendix B Proofs of the technical results
B.1 Proof of Proposition 3
We give Lemma 2.1 of Zhivotovskiy, (2024) or an extension of Proposition 2.1 of Catoni and Giulini, (2017) to cases with unbounded loss functions for the sake of completeness.
Lemma 7.
Fix a probability measure $\mu$ on $(\mathbb{R}^{p},\mathcal{B}(\mathbb{R}^{p}))$ and a measurable function $\phi:\mathbb{M}\times\mathbb{R}^{p}\to\mathbb{R}$. Assume that (i) $\int_{\mathbb{R}^{p}}\int_{\mathbb{M}}\exp(\phi(v,\boldsymbol{w}))P^{V}(\mathrm{d}v)\mu(\mathrm{d}\boldsymbol{w})<\infty$ and (ii) $|\phi(v,\boldsymbol{w})|<\infty$ $P^{V}\otimes\mu$-almost everywhere. For any $\delta\in(0,1]$, with probability at least $1-\delta$, for any probability measure $\rho$ on $(\mathbb{R}^{p},\mathcal{B}(\mathbb{R}^{p}))$,
$\sum_{i=1}^{n}\int_{\mathbb{R}^{p}}\phi\left(V_{i},\boldsymbol{w}\right)\rho\left(\mathrm{d}\boldsymbol{w}\right)\leq n\int_{\mathbb{R}^{p}}\log\left(\int_{\mathbb{M}}\exp\left(\phi\left(v,\boldsymbol{w}\right)\right)P^{V}(\mathrm{d}v)\right)\rho\left(\mathrm{d}\boldsymbol{w}\right)+H\left(\rho\|\mu\right)+\log\delta^{-1}.$ (34)
Proof.
The proof is identical to that of Proposition 2.1 of Catoni and Giulini, (2017) except for the unboundedness of integrands. We obtain the following fact via a variational representation (see Section 5.2 of Catoni,, 2004): for any measurable function $h:(\mathbb{R}^{p},\mathcal{B}(\mathbb{R}^{p}))\to(\mathbb{R},\mathcal{B}(\mathbb{R}))$,
$\sup_{b\in\mathbb{R}}\log\left(\int_{\mathbb{R}^{p}}\exp\left(\min\left\{b,h\left(\boldsymbol{w}\right)\right\}\right)\mu\left(\mathrm{d}\boldsymbol{w}\right)\right)=\sup_{\rho}\left(\sup_{b\in\mathbb{R}}\int_{\mathbb{R}^{p}}\min\left\{b,h\left(\boldsymbol{w}\right)\right\}\rho\left(\mathrm{d}\boldsymbol{w}\right)-H\left(\rho\|\mu\right)\right),$ (35)
where the supremum in $\rho$ is taken over all the probability measures on $(\mathbb{R}^{p},\mathcal{B}(\mathbb{R}^{p}))$ and both the left-hand side and the right-hand side can be infinite (we let $\infty$-$\infty$=$\infty$ - ∞ = - ∞ as Catoni,, 2004). Fubini’s theorem and Condition (i) give that
$\mathbb{E}_{V}\exp\sup_{\rho}\left(\mathbb{E}_{\boldsymbol{W}\sim\rho}\left(\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right)-H\left(\rho\|\mu\right)\right)$
$=\mathbb{E}_{V}\exp\sup_{\rho}\left(\sup_{b\in\mathbb{R}}\mathbb{E}_{\boldsymbol{W}\sim\rho}\min\left\{b,\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right\}-H\left(\rho\|\mu\right)\right)$
$=\mathbb{E}_{V}\exp\sup_{b\in\mathbb{R}}\log\mathbb{E}_{\boldsymbol{W}\sim\mu}\exp\min\left\{b,\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right\}$
$=\mathbb{E}_{V}\sup_{b\in\mathbb{R}}\mathbb{E}_{\boldsymbol{W}\sim\mu}\exp\min\left\{b,\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right\}$
$=\mathbb{E}_{V}\mathbb{E}_{\boldsymbol{W}\sim\mu}\exp\left(\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right)$
$=\mathbb{E}_{\boldsymbol{W}\sim\mu}\mathbb{E}_{V}\exp\left(\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right)$
$=\mathbb{E}_{\boldsymbol{W}\sim\mu}\prod_{i=1}^{n}\frac{\mathbb{E}_{V}\exp\phi\left(V_{i},\boldsymbol{W}\right)}{\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)}$
$=1;$
the denominator does not vanish $\mu$-almost surely owing to Condition (ii). Therefore, by setting a random variable $\xi$ such that
$\xi:=\sup_{\rho}\left(\mathbb{E}_{\boldsymbol{W}\sim\rho}\left(\sum_{i=1}^{n}\phi\left(V_{i},\boldsymbol{W}\right)-n\log\left(\mathbb{E}_{V}\exp\phi\left(V_{1},\boldsymbol{W}\right)\right)\right)-H\left(\rho\|\mu\right)\right),$
Markov’s inequality yields that $\mathbb{P}(\xi\geq\log\delta^{-1})\leq\delta\mathbb{E}\exp(\xi)=\delta$; this is the desired conclusion. ∎
Proof of Proposition 3.
We give only the first statement since the second one is quite parallel by flipping the sign of $f$. We also consider only the case where $C_{t}>0$; otherwise the proof is trivial. Itô’s formula gives that for all $\boldsymbol{\theta}\in B[R]$,
$\mathbb{E}_{\boldsymbol{W}}\left[f\left(V_{i},\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]=f(V_{i},\boldsymbol{\theta})+\frac{1}{2}\int_{0}^{t}\mathbb{E}_{\boldsymbol{W}}\left[\left.\Delta_{\boldsymbol{w}}f\left(V_{i},\boldsymbol{w}\right)\right|_{\boldsymbol{w}=\boldsymbol{W}_{t}^{\boldsymbol{\theta}}}\right]\mathrm{d}t$ (36)
almost surely since the local martingale term is martingale by (i)-(a), and the interchange of expectations holds by (i)-(b). We have
$H\left(\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\left\|\boldsymbol{W}_{t}^{\mathbf{0}}\right.\right)=\frac{1}{2t}\left\|\boldsymbol{\theta}\right\|^{2}\leq\frac{R^{2}}{2t}$ (37)
(we use the law of $\boldsymbol{W}_{t}^{\mathbf{0}}=\boldsymbol{W}_{t}$ as the “prior” and those of $\boldsymbol{W}_{t}^{\boldsymbol{\theta}}$ with $\boldsymbol{\theta}\in B[R]$ as the “posteriors”). By Lemma 7 (Condition (ii) of Lemma 7 is satisfied by Condition (ii) of the proposition), for any $t>0$ and $\lambda>0$ such that $\mathbb{E}_{V,\boldsymbol{W}}[\exp(\lambda(f(V_{1},\boldsymbol{W}_{t})-\mathbb{E}_{V}[f(V_{1},\boldsymbol{W}_{t})]))]<\infty$, with probability at least $1-\delta$, for all $\boldsymbol{\theta}\in B[R]$,
$\lambda\sum_{i=1}^{n}\left(\mathbb{E}_{V,\boldsymbol{W}}\left[f\left(V_{i},\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]-\mathbb{E}_{\boldsymbol{W}}\left[f\left(V_{i},\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right]\right)$ (38)
$\leq n\mathbb{E}_{\boldsymbol{W}}\log\mathbb{E}_{V}\exp\left(\lambda\left(f\left(V_{1},\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)-\mathbb{E}_{V}f\left(V_{1},\boldsymbol{W}_{t}^{\boldsymbol{\theta}}\right)\right)\right)+\frac{R^{2}}{2t}+\log\delta^{-1}$
$\leq\frac{n\lambda^{2}C_{t}}{8}+\frac{R^{2}}{2t}+\log\delta^{-1}$ (39)
Combining this inequality with Eq. (36), dividing by $\lambda n$, and fixing
$\lambda=\sqrt{\frac{8(R^{2}/(2t)+\log\delta^{-1})}{nC_{t}}},$
we obtain the conclusion. ∎
B.2 Proof of Proposition 4
We prove several preliminary lemmas for Proposition 4.
Lemma 8.
Suppose that Assumption 1 holds. It holds that for all $t>0$, $\boldsymbol{\theta}\in B[R]$, and $\boldsymbol{z}=(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})\in\mathbb{R}^{np}$,
$\left\|\nabla_{\boldsymbol{z}}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z})\right\|\leq\sqrt{\frac{2t\left\|\Sigma\right\|}{\pi n}}\left(4+\sqrt{\frac{R^{2}}{16n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\right).$ (40)
Proof.
In this proof, we frequently use the notation $\zeta\sim\mathcal{N}(0,1)$, $\mu_{i}=\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{z}_{i}\rangle$, and $\lambda_{s,i}=\sqrt{s\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\rangle}$ for all $s\in[0,t]$, and the following fact:
$\left[\begin{matrix}\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\\ \left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{u}_{i}\right\rangle\end{matrix}\right]\sim\mathcal{N}\left(\left[\begin{matrix}\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{z}_{i}\right\rangle\\ \left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}\right\rangle\end{matrix}\right],s\left[\begin{matrix}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle&\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\\ \left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle&\left\langle\Lambda\boldsymbol{u}_{i},\boldsymbol{u}_{i}\right\rangle\end{matrix}\right]\right).$ (41)
We choose arbitrary $\boldsymbol{u}=(\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{n})\in\mathbb{S}^{np-1}$ and give a uniform bound for
$\left\langle\nabla_{\boldsymbol{z}}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z}),\boldsymbol{u}\right\rangle$ $=\frac{1}{n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\right]\mathrm{d}s$
$\quad+\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{u}_{i}\right\rangle\right]\mathrm{d}s$
$=\frac{1}{n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\right]\mathrm{d}s$
$\quad+\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\right]\mathrm{d}s$
$\quad+\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle-\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{z}_{i}\right\rangle\right)\right]\mathrm{d}s$
$=\frac{1}{n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\sigma\left(1-\sigma\right)\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\mathrm{d}s$
$\quad+\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\mathrm{d}s$
$\quad+\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\lambda_{s,i}\zeta\right]\mathrm{d}s;$ (42)
the interchange of limiting operations in the first equality holds since $\|(\sigma(1-\sigma))^{\prime}\|_{\infty}\|\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}}\|$ dominates the integrand, and the second equality holds since for any $s\in(0,t]$,
$\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{u}_{i}\right\rangle\right]$
$=\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\left(\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}\right\rangle+\frac{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle}{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle-\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{z}_{i}\right\rangle\right)\right)\right].$
We use the following fact repetitively: for any $\{a_{1},\ldots,a_{n}\}\subset\mathbb{R}^{n}$, $\{b_{1},\ldots,b_{n}\}\subset[0,\infty)^{n}$ with $|a_{i}|\leq b_{i}$, $\boldsymbol{v}=(\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{n})\in\mathbb{R}^{np}$ and $\boldsymbol{u}=(\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{n})\in\mathbb{S}^{np-1}$, by choosing $\bar{\boldsymbol{u}}=(\bar{\boldsymbol{u}}_{1},\ldots,\bar{\boldsymbol{u}}_{n})=(\mathrm{sgn}(\langle\boldsymbol{v}_{1},\boldsymbol{u}_{1}\rangle)\boldsymbol{u}_{1},\ldots,\mathrm{sgn}(\langle\boldsymbol{v}_{n},\boldsymbol{u}_{n}\rangle)\boldsymbol{u}_{n})\in\mathbb{S}^{np-1}$,
$\sum_{i=1}^{n}a_{i}\langle\boldsymbol{u}_{i},\boldsymbol{v}_{i}\rangle\leq\sum_{i=1}^{n}|a_{i}\langle\boldsymbol{u}_{i},\boldsymbol{v}_{i}\rangle|\leq\sum_{i=1}^{n}b_{i}\langle\bar{\boldsymbol{u}}_{i},\boldsymbol{v}_{i}\rangle\leq\sup_{\tilde{\boldsymbol{u}}\in\mathbb{S}^{np-1}}\sum_{i=1}^{n}b_{i}\langle\tilde{\boldsymbol{u}}_{i},\boldsymbol{v}_{i}\rangle=\sqrt{\sum_{i=1}^{n}b_{i}^{2}\|\boldsymbol{v}_{i}\|^{2}}.$ (43)
(Step 1) We give a bound for the first term on the most right-hand side of (42). Stein’s lemma. the fact $\sigma^{\prime}=\sigma(1-\sigma)$, Hölder’s inequality, $\|\sigma\|_{\infty}\leq 1$, and $\mathbb{E}[|\zeta|]=\sqrt{2/\pi}$ give that
$\left|\mathbb{E}\left[\sigma\left(1-\sigma\right)\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|$ $=\left|\mathbb{E}\left[\sigma^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\partial_{\zeta}\sigma\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|$
$=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\sigma\left(\mu_{i}+\lambda_{s,i}\zeta\right)\zeta\right]\right|\leq\frac{1}{\lambda_{s,i}}\sqrt{\frac{2}{\pi}}.$
Therefore, for some $\boldsymbol{u}^{\prime}=(\boldsymbol{u}_{1}^{\prime},\ldots,\boldsymbol{u}_{n}^{\prime})\in\mathbb{S}^{np-1}$,
$\frac{1}{n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\sigma\left(1-\sigma\right)\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\mathrm{d}s$ $=\frac{1}{n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\frac{1}{\lambda_{s,i}}\mathbb{E}\left[\sigma\left(\mu_{i}+\lambda_{s,i}\zeta\right)\zeta\right]\mathrm{d}s$
$\leq\frac{\sqrt{2/\pi}}{n}\sum_{i=1}^{n}\mathrm{d}s\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}^{\prime}\right\rangle\int_{0}^{t}\frac{1}{\sqrt{s\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\rangle}}\mathrm{d}s$
$\leq\frac{2\sqrt{2/\pi}}{n}\sum_{i=1}^{n}\frac{\sqrt{t}}{\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\mathrm{d}s\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}^{\prime}\right\rangle$
$\leq 2\sqrt{\frac{2t}{\pi n}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left\|\frac{1}{\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\Lambda\boldsymbol{z}_{i}\right\|^{2}}$
$\leq 2\sqrt{\frac{2t\left\|\Sigma\right\|}{\pi n}},$
where the second inequality is by the fact (43).
(Step 2) In the next place, we give an upper bound for the second term on the most right-hand side of (42). Stein’s lemma yields that
$\left|\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|$ $=\left|\frac{1}{\lambda_{s,i}}\mathbb{E}\left[\lambda_{s,i}\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|$
$=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\partial_{\zeta}\left(\sigma\left(1-\sigma\right)\right)\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\right|$
$=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)\left(\mu_{i}+\lambda_{s,i}\zeta\right)\zeta\right]\right|$
$\leq\frac{\sqrt{2/\pi}}{4\lambda_{s,i}},$
where the last inequality uses Hölder’s inequality and $\|\sigma(1-\sigma)\|_{\infty}\leq 1/4$. For some $\boldsymbol{u}^{\prime\prime}=(\boldsymbol{u}_{1}^{\prime\prime},\ldots,\boldsymbol{u}_{n}^{\prime\prime})\in\mathbb{S}^{np-1}$, by the fact (43),
$\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right]\mathrm{d}s$
$\leq\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}^{\prime\prime}\right\rangle\int_{0}^{t}\frac{\sqrt{2/\pi}}{4\sqrt{s\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\mathrm{d}s$
$\leq\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\frac{\sqrt{2t/\pi}}{2\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\left\langle\boldsymbol{\theta},U\Lambda^{1/2}\boldsymbol{u}_{i}^{\prime\prime}\right\rangle$
$=\frac{\sqrt{2t/\pi}}{4n}\sum_{i=1}^{n}\left\langle\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\Lambda^{1/2}U^{\top}\boldsymbol{\theta},\boldsymbol{u}_{i}^{\prime\prime}\right\rangle$
$\leq\frac{\sqrt{2t/\pi}}{4n}\left\|\left(\sqrt{\left\langle\Lambda\boldsymbol{z}_{1},\boldsymbol{z}_{1}\right\rangle}\Lambda^{1/2}U^{\top}\boldsymbol{\theta},\ldots,\sqrt{\left\langle\Lambda\boldsymbol{z}_{n},\boldsymbol{z}_{n}\right\rangle}\Lambda^{1/2}U^{\top}\boldsymbol{\theta}\right)\right\|$
$=\frac{\sqrt{2t/\pi}}{4n}\sqrt{\sum_{i=1}^{n}\left\|\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\Lambda^{1/2}U^{\top}\boldsymbol{\theta}\right\|^{2}}$
$=\frac{\sqrt{2t/\pi}}{4n}\sqrt{\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\left\|\Lambda^{1/2}U^{\top}\boldsymbol{\theta}\right\|^{2}}$
$\leq\sqrt{\frac{2t\left\|\Sigma\right\|}{\pi n}}\sqrt{\frac{R^{2}}{16n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}.$
(Step 3) We examine the third term on the most right-hand side of (42). It holds that
$\left|\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\lambda_{s,i}\zeta\right]\right|$ $=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\left(\partial_{\zeta}^{2}\sigma\left(\mu_{i}+\lambda_{s,i}\zeta\right)\right)\zeta\right]\right|$
$=\frac{1}{\lambda_{s,i}}\left|\mathbb{E}\left[\sigma\left(\mu_{i}+\lambda_{s,i}\zeta\right)\left(\zeta^{3}-3\zeta\right)\right]\right|$
$\leq\frac{1}{\lambda_{s,i}}\mathbb{E}\left[\left|\zeta^{3}-3\zeta\right|\right]$
$\leq\frac{2\sqrt{2/\pi}}{\lambda_{s,i}},$
where the second equality is by Lemma 6-(ii) with $d=1$, the third line is by Hölder’s inequality, and the last inequality holds by $\mathbb{E}[|\zeta^{3}-3\zeta|]=(1+4e^{-3/2})\sqrt{2/\pi}<2\sqrt{2/\pi}$. By the fact (43), for some $\boldsymbol{u}^{\prime\prime\prime}=(\boldsymbol{u}_{1}^{\prime\prime\prime},\ldots,\boldsymbol{u}_{n}^{\prime\prime\prime})\in\mathbb{S}^{np-1}$,
$\sum_{i=1}^{n}\frac{1}{2n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}\right\rangle\int_{0}^{t}\mathbb{E}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\mu_{i}+\lambda_{s,i}\zeta\right)\lambda_{s,i}\zeta\right]\mathrm{d}s$ $\leq 2\sum_{i=1}^{n}\frac{\sqrt{2t/\pi}}{n}\left\langle\frac{1}{\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\Lambda\boldsymbol{z}_{i},\boldsymbol{u}_{i}^{\prime\prime\prime}\right\rangle$
$=\frac{2\sqrt{2t/\pi}}{n}\sqrt{\sum_{i=1}^{n}\left\|\frac{1}{\sqrt{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}}\Lambda\boldsymbol{z}_{i}\right\|^{2}}$
$\leq\frac{2\sqrt{2t/\pi}}{n}\sqrt{\sum_{i=1}^{n}\left(\frac{1}{\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\left\|\Lambda\boldsymbol{z}_{i}\right\|^{2}\right)}$
$\leq 2\sqrt{\frac{2t\|\Sigma\|}{\pi n}}.$
(Step 4) Therefore, for all $\boldsymbol{u}\in\mathbb{S}^{np-1}$,
$\left\langle\nabla_{\boldsymbol{z}}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z}),\boldsymbol{u}\right\rangle\leq\sqrt{\frac{2t\left\|\Sigma\right\|}{\pi n}}\left(4+\sqrt{\frac{R^{2}}{16n}\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}\right).$ (44)
This is the desired conclusion. ∎
The next lemma is a version of Lemmas 3.1 and 3.2 of Adamczak, (2015).
Lemma 9.
Let $\{S_{v};v>0\}$ be a sequence of real-valued integrable random variables, $Z$ be a real-valued integrable random variable, and $a,b>0$ such that for all $v>0$ and $u>0$,
$\mathbb{P}\left(|S_{v}-\mathbb{E}[S_{v}]|\geq s\right)\leq 2\exp\left(-\frac{u^{2}}{a+\sqrt{bv}}\right)$
and
$\mathbb{P}\left(S_{v}\neq Z\right)\leq 2\exp\left(-v/b\right).$
Then for any $v>0$,
$\mathbb{P}\left(|Z-\mathbb{E}[Z]|\geq v\right)\leq 2\exp\left(-\frac{1}{6}\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\right).$ (45)
Proof.
Let $\mathrm{Med}(Z)$ denote the median of $Z$. By the proof of Lemma 3.1 of Adamczak, (2015), we know that for any $v>0$,
$\mathbb{P}\left(|Z-\mathrm{Med}(Z)|\geq v\right)\leq 27\exp\left(-\min\left\{\frac{v^{2}}{4a^{2}},\frac{v}{b}\right\}\right).$ (46)
We also obtain that
$\left|\mathbb{E}[Z]-\mathrm{Med}(Z)\right|\leq\mathbb{E}\left[\left|Z-\mathrm{Med}(Z)\right|\right]\leq 27\int_{0}^{\infty}\exp\left(-\min\left\{\frac{v^{2}}{4a^{2}},\frac{v}{b}\right\}\right)\mathrm{d}v\leq 27(\sqrt{\pi}a+b).$
For $v>(22/21)(27(\sqrt{\pi}a+b))$, we have
$\mathbb{P}\left(\left|Z-\mathbb{E}[Z]\right|\geq v\right)\leq\mathbb{P}\left(\left|Z-\mathrm{Med}(Z)\right|\geq v/22\right)$ $\leq\min\left\{1,27\exp\left(-\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\right)\right\}$
$\leq\min\left\{1,2\exp\left(-\frac{1}{5}\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\right)\right\}.$
If $v\leq(22/21)(27(\sqrt{\pi}a+b))$, then it is sufficient to see that
$\frac{1}{\log 2}\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\leq 6.$ (47)
If $a\geq b$,
$\frac{1}{\log 2}\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\leq\frac{1}{\log 2}\frac{(22/21)^{2}(27(\sqrt{\pi}a+b))^{2}}{(44a)^{2}}=\frac{(27)^{2}(\sqrt{\pi}+b/a)^{2}}{(42)^{2}\log 2}<5,$
and if $b\geq a$,
$\frac{1}{\log 2}\min\left\{\frac{v^{2}}{(44a)^{2}},\frac{v}{22b}\right\}\leq\frac{1}{\log 2}\frac{27(\sqrt{\pi}a+b)}{21b}\leq\frac{27(\sqrt{\pi}+1)}{21\log 2}<6.$
We obtain the conclusion. ∎
Proof of Proposition 4.
We only prove the first statement. The proof is partially adopted from Adamczak, (2015). We have that for any $\boldsymbol{z},\boldsymbol{z}^{\prime}\in\mathbb{R}^{np}$,
$\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z})-\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z}^{\prime})\leq\sqrt{\frac{2t\left\|\Sigma\right\|}{\pi n}}\left(4+\sqrt{\frac{R^{2}}{16n}\left(\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\right)\vee\left(\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i}^{\prime},\boldsymbol{z}_{i}^{\prime}\right\rangle\right)}\right)\left\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\right\|.$ (48)
We fix a positive number $v>0$, a convex set
$B:=\left\{\boldsymbol{z}\in\mathbb{R}^{np};\sqrt{\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle}<\mathbb{E}\left[\sqrt{\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle}\right]+\sqrt{v\|\Sigma\|}\right\}\subset\mathbb{R}^{np}$ (49)
and $((16\sqrt{n}/R+\mathbb{E}[\sqrt{(\sum_{i=1}^{n}\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\rangle)}])/\sqrt{\|\Sigma\|}+\sqrt{v})$-Lipschitz function $\varphi:\mathbb{R}^{np}\to\mathbb{R}$ such that for all $\boldsymbol{z}\in B$,
$\varphi(\boldsymbol{z})=\sqrt{\frac{16\pi n^{2}}{2tR^{2}\|\Sigma\|^{2}}}\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{z}).$ (50)
Then by the concentration property and the Cauchy–Schwarz inequality, for all $u>0$,
$\mathbb{P}\left(\left|\varphi(\boldsymbol{Z})-\mathbb{E}[\varphi(\boldsymbol{Z})]\right|\geq u\right)$ $\leq 2\exp\left(-\frac{u^{2}}{K^{2}\left(\left(16\sqrt{n}/R+\mathbb{E}\left[\sqrt{(\sum_{i=1}^{n}\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\rangle)}\right]\right)/\sqrt{\|\Sigma\|}+\sqrt{v}\right)^{2}}\right)$
$\leq 2\exp\left(-\frac{u^{2}}{K^{2}\left(\left(16\sqrt{n}/R+\sqrt{n\Sigma}\right)/\sqrt{\|\Sigma\|}+\sqrt{v}\right)^{2}}\right).$ (51)
The concentration property and the $1$-Lipschitz continuity of $\sqrt{\sum_{i=1}^{n}\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\rangle}/\sqrt{\|\Sigma\|}$ also yield that for all $v>0$,
$\mathbb{P}\left(\sqrt{\left(\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle\right)}\geq\mathbb{E}\left[\sqrt{\left(\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle\right)}\right]+\sqrt{v\|\Sigma\|}\right)\leq 2\exp\left(-\frac{v}{K^{2}}\right).$ (52)
Hence by Lemma 9 with $a=K((16\sqrt{n}/R)+\sqrt{n\mathrm{tr}(\Sigma)})/\sqrt{\|\Sigma\|}$ and $b=K^{2}$, for all $u>0$,
$\mathbb{P}\left(\sqrt{\frac{16\pi n^{2}}{2tR^{2}\|\Sigma\|^{2}}}\left|\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})-\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})\right]\right|\geq u\right)$
$\leq 2\exp\left(-\frac{1}{6K^{2}}\min\left\{\frac{u^{2}}{44^{2}\left(16\sqrt{n}/R+\sqrt{n\mathrm{tr}(\Sigma)}\right)^{2}/\|\Sigma\|^{2}},\frac{u}{22}\right\}\right).$
Therefore,
$\mathbb{P}\left(\left|\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})-\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}(\boldsymbol{Z})\right]\right|\geq u\right)$
$\leq 2\exp\left(-\frac{1}{6K^{2}}\min\left\{\frac{n^{2}u^{2}}{11^{2}R^{2}(2t/\pi)\left(\frac{16\sqrt{n}}{R}+\sqrt{n\mathrm{tr}(\Sigma)}\right)^{2}},\frac{2nu}{11R\sqrt{2t/\pi}\|\Sigma\|}\right\}\right)$
$\leq 2\exp\left(-\frac{n}{6K^{2}}\min\left\{\frac{u^{2}}{121(2t/\pi)\left(16+R\sqrt{\mathrm{tr}(\Sigma)}\right)^{2}},\frac{u}{5R\sqrt{t}\|\Sigma\|}\right\}\right)$
$\leq 2\exp\left(-\frac{n}{6K^{2}}\min\left\{\frac{u^{2}}{155t\left(256+R^{2}\mathrm{tr}\left(\Sigma\right)\right)},\frac{u}{5\sqrt{t}R\|\Sigma\|}\right\}\right).$
Hence we obtain the conclusion. ∎
B.3 Proof of Proposition 5
Lemma 10.
For arbitrary function $f\in\mathcal{C}_{b}^{2}(\mathbb{R};\mathbb{R})$ and $t\in[0,1]$, we have
$\int_{0}^{t}\frac{1}{s}\int_{-\infty}^{\infty}f\left(\sqrt{s}x\right)\left(x^{2}-1\right)\frac{\exp(-x^{2}/2)}{\sqrt{2\pi}}\mathrm{d}x\mathrm{d}s=2\left(\int_{-\infty}^{\infty}f(\sqrt{t}x)\frac{\exp(-x^{2}/2)}{\sqrt{2\pi}}\mathrm{d}x-f(0)\right).$ (53)
Proof.
We change variables $s=e^{-2r}$ ($\mathrm{d}s=-2e^{-2r}\mathrm{d}r=-2s\mathrm{d}r$) and then for $T_{0}:=\log(1/\sqrt{t})$,
$\int_{0}^{t}\frac{1}{s}\int_{-\infty}^{\infty}f\left(\sqrt{s}x\right)\left(x^{2}-1\right)\frac{\exp\left(-x^{2}/2\right)}{\sqrt{2\pi}}\mathrm{d}x\mathrm{d}s=2\int_{T_{0}}^{\infty}\int_{-\infty}^{\infty}f(e^{-r}x)\left(x^{2}-1\right)\frac{\exp\left(-x^{2}/2\right)}{\sqrt{2\pi}}\mathrm{d}x\mathrm{d}r.$ (54)
Let us consider a $1$-dimensional stationary Ornstein–Uhlenbeck process $\{X_{t};t\geq 0\}$ defined by the following stochastic differential equation:
$\mathrm{d}X_{t}=-X_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\ X_{0}\sim\mathcal{N}(0,1),$ (55)
where $\{W_{t};t\geq 0\}$ is a $1$-dimensional standard Brownian motion independent of $X_{0}$. Note that for each $t\geq 0$, $X_{t}\sim\mathcal{N}(0,1)$.
For any $T_{1}\geq T_{0}$,
$\mathbb{E}\left[f\left(e^{-T_{1}}X_{T_{1}}\right)\right]$ $=\mathbb{E}\left[f\left(e^{-T_{0}}X_{T_{0}}\right)\right]+\int_{T_{0}}^{T_{1}}\mathbb{E}\left[\left.-X_{s}\partial_{x}f(e^{-s}x)+\partial_{s}f(e^{-s}x)+\partial_{x}^{2}f(e^{-s}x)\right|_{x=X_{s}}\right]\mathrm{d}s$
$=\mathbb{E}\left[f\left(e^{-T_{0}}X_{T_{0}}\right)\right]+\int_{T_{0}}^{T_{1}}\mathbb{E}\left[\left.-2X_{s}\partial_{x}f(e^{-s}x)\right|_{x=X_{s}}+f(e^{-s}X_{s})(X_{s}^{2}-1)\right]\mathrm{d}s$
$=\mathbb{E}\left[f\left(e^{-T_{0}}X_{T_{0}}\right)\right]-\int_{T_{0}}^{T_{1}}\mathbb{E}\left[f(e^{-s}X_{s})(X_{s}^{2}-1)\right]\mathrm{d}s,$
by Itô’s formula, Lemma 6-(ii) with $d=0$, and the fact that $\partial_{s}f(e^{-s}x)=-e^{-s}xf^{\prime}(e^{-s}x)=-x\partial_{x}f(e^{-s}x)$. Taking the limit $T_{1}\to\infty$, we obtain
$\int_{T_{0}}^{\infty}\mathbb{E}\left[f(e^{-s}X_{s})(X_{s}^{2}-1)\right]\mathrm{d}s=\mathbb{E}\left[f\left(e^{-T_{0}}X_{T_{0}}\right)\right]-f(0).$ (56)
This left-hand side is identical to the right-hand side of Eq. (54) up to constant $2$. ∎
We show a concise version of Lemma 5 of Meir and Zhang, (2003) for self-containedness; see also Theorem 4.12 of Ledoux and Talagrand, (1991) and Proposition 4.3 of Bach, (2023).
Lemma 11.
For arbitrary sequences of functions $\{a_{i}\}$ with $a_{i}:B[R]\to\mathbb{R}$, a function $b:B[R]\to\mathbb{R}$, and a sequence of $L_{i}$-Lipschitz functions $\{\varphi_{i}\}$ with $\varphi_{i}:\mathbb{R}\to\mathbb{R}$,
$\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{n}\varepsilon_{i}\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)\right)\right]\leq\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{n}\varepsilon_{i}L_{i}a_{i}\left(\boldsymbol{\theta}\right)\right)\right],$ (57)
where $\{\varepsilon_{i};i=1,\ldots,n\}$ is an i.i.d. sequence of Rademacher random variables.
Proof.
We prove the statement by induction as Meir and Zhang, (2003). We consider $n=0$, and then the statement immediately holds. Let us assume that the statement is true for some $n=m-1$ with $m\geq 2$. It holds that
$\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m}\varepsilon_{i}\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)\right)\right]$
$=\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)+\varphi_{m}\left(a_{m}\left(\boldsymbol{\theta}\right)\right)\right)\right]$
$\quad+\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)-\varphi_{m}\left(a_{m}\left(\boldsymbol{\theta}\right)\right)\right)\right]$
$=\mathbb{E}\left[\sup_{\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in B[R]}\left(\frac{b\left(\boldsymbol{\theta}\right)+b\left(\boldsymbol{\theta}^{\prime}\right)}{2}+\sum_{i=1}^{m-1}\varepsilon_{i}\frac{\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)+\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}^{\prime}\right)\right)}{2}+\frac{\varphi_{m}\left(a_{m}\left(\boldsymbol{\theta}\right)\right)-\varphi_{m}\left(a_{m}\left(\boldsymbol{\theta}^{\prime}\right)\right)}{2}\right)\right]$
$\leq\mathbb{E}\left[\sup_{\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in B[R]}\left(\frac{b\left(\boldsymbol{\theta}\right)+b\left(\boldsymbol{\theta}^{\prime}\right)}{2}+\sum_{i=1}^{m-1}\varepsilon_{i}\frac{\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)+\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}^{\prime}\right)\right)}{2}+\frac{L_{m}\left|a_{m}\left(\boldsymbol{\theta}\right)-a_{m}\left(\boldsymbol{\theta}^{\prime}\right)\right|}{2}\right)\right]$
$=\mathbb{E}\left[\sup_{\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in B[R]}\left(\frac{b\left(\boldsymbol{\theta}\right)+b\left(\boldsymbol{\theta}^{\prime}\right)}{2}+\sum_{i=1}^{m-1}\varepsilon_{i}\frac{\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}\right)\right)+\varphi_{i}\left(a_{i}\left(\boldsymbol{\theta}^{\prime}\right)\right)}{2}+\frac{L_{m}\left(a_{m}\left(\boldsymbol{\theta}\right)-a_{m}\left(\boldsymbol{\theta}^{\prime}\right)\right)}{2}\right)\right]$
$=\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}\varphi_{i}\left(a\left(\boldsymbol{\theta}\right)\right)+L_{m}a_{m}\left(\boldsymbol{\theta}\right)\right)\right]$
$\quad+\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}\varphi_{i}\left(a\left(\boldsymbol{\theta}\right)\right)-L_{m}a_{m}\left(\boldsymbol{\theta}\right)\right)\right]$
$\leq\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}L_{i}a_{i}\left(\boldsymbol{\theta}\right)+L_{m}a_{m}\left(\boldsymbol{\theta}\right)\right)\right]$
$\quad+\frac{1}{2}\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m-1}\varepsilon_{i}L_{i}a_{i}\left(\boldsymbol{\theta}\right)-L_{m}a_{m}\left(\boldsymbol{\theta}\right)\right)\right]$
$=\mathbb{E}\left[\sup_{\boldsymbol{\theta}\in B[R]}\left(b\left(\boldsymbol{\theta}\right)+\sum_{i=1}^{m}\varepsilon_{i}L_{i}a_{i}\left(\boldsymbol{\theta}\right)\right)\right],$
where the second last inequality is given by the induction hypothesis. ∎
Proof of Proposition 5.
We only give proof of the first statement as the second statement is parallel. Proposition 4.2 of Bach, (2023) yields that for each $t\in(0,1]$,
$\mathbb{E}_{\boldsymbol{Z}}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}\left(\boldsymbol{Z}\right)\right]\leq 2\mathbb{E}_{\boldsymbol{\varepsilon},\boldsymbol{Z}}\left[\sup_{\boldsymbol{\theta}\in B[R]}\frac{1}{2n}\sum_{i=1}^{n}\varepsilon_{i}\int_{0}^{t}\mathbb{E}_{\boldsymbol{W}}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{Z}_{i}\right\rangle\right)\right]\mathrm{d}s\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle\right],$ (58)
where $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{n})$ is an $\mathbb{R}^{n}$-valued random vector whose components are independent Rademacher random variables, which is also independent of $\boldsymbol{W}$ and $\boldsymbol{Z}$. We set $\lambda_{i}=\lambda_{i}(\boldsymbol{z}_{i})=\sqrt{\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\rangle},\nu_{i}=\nu_{i}(\boldsymbol{z}_{i},\boldsymbol{\theta})=\lambda_{i}^{-1}\langle U\Lambda^{1/2}\boldsymbol{z}_{i},\boldsymbol{\theta}\rangle$, and $\zeta\sim\mathcal{N}(0,1)$; then $\mathbb{E}_{\boldsymbol{W}}[\sigma(1-\sigma)(\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{Z}_{i}\rangle)]=\mathbb{E}_{\zeta}[\sigma(1-\sigma)(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta))].$ By Lemma 6, we have that for any $s>0$ and $i=1,\ldots,n$,
$\partial_{\nu_{i}}\mathbb{E}_{\zeta}\left[\sigma\left(1-\sigma\right)\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]$ $=\mathbb{E}_{\zeta}\left[\left(\sigma\left(1-\sigma\right)\right)^{\prime}\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]\lambda_{i}$
$=\mathbb{E}_{\zeta}\left[(\partial_{\zeta}^{2}\sigma)\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]\frac{\lambda_{i}}{s\lambda_{i}^{2}}$
$=\mathbb{E}_{\zeta}\left[\sigma\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\left(\zeta^{2}-1\right)\right]\frac{1}{s\lambda_{i}};$
the interchange of limiting operations in the first line holds as the partial derivative of the integrand on the left-hand side is dominated by $\lambda_{i}\|(\sigma(1-\sigma))^{\prime}\|_{\infty}<\infty$ on $\mathbb{R}$. The left-hand side itself is also dominated by $\lambda_{i}\|(\sigma(1-\sigma))^{\prime}\|_{\infty}$ on $(0,t]$. Therefore, Lemma 10 gives that for any $t\in(0,1]$,
$\left|\partial_{\nu_{i}}\int_{0}^{t}\mathbb{E}_{\zeta}\left[\sigma\left(1-\sigma\right)\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]\mathrm{d}s\right|$ $=\left|\int_{0}^{t}\partial_{\nu_{i}}\mathbb{E}_{\zeta}\left[\sigma\left(1-\sigma\right)\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]\mathrm{d}s\right|$
$=\left|\int_{0}^{t}\mathbb{E}_{\zeta}\left[\sigma\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\left(\zeta^{2}-1\right)\right]\frac{1}{s\lambda_{i}}\mathrm{d}s\right|$
$=\frac{2}{\lambda_{i}}\left|\mathbb{E}\left[\sigma(\lambda_{i}(\nu_{i}+\sqrt{t}\zeta))\right]-\sigma(\lambda_{i}(\nu_{i}))\right|$
$\leq\frac{2}{\lambda_{i}}.$
Therefore, for any $t\in(0,1]$,
$\left|\partial_{\nu_{i}}\int_{0}^{t}\mathbb{E}_{\zeta}\left[\sigma\left(1-\sigma\right)\left(\lambda_{i}(\nu_{i}+\sqrt{s}\zeta)\right)\right]\mathrm{d}s\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\right|\leq 2\lambda_{i}.$ (59)
By Lemma 11 (with $a_{i}=\nu_{i}$, $b=0$, and $L_{i}=2\lambda_{i}$), we obtain that for all $t\in(0,1]$ and $\boldsymbol{z}\in\mathbb{R}^{np}$,
$\mathbb{E}_{\boldsymbol{\varepsilon}}\left[\sup_{\boldsymbol{\theta}\in B[R]}\sum_{i=1}^{n}\varepsilon_{i}\int_{0}^{t}\mathbb{E}_{\boldsymbol{W}}\left[\sigma\left(1-\sigma\right)\left(\left\langle\tilde{\boldsymbol{W}}_{s}^{\boldsymbol{\theta}},\boldsymbol{z}_{i}\right\rangle\right)\right]\mathrm{d}s\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\right]$ $\leq 2\mathbb{E}_{\boldsymbol{\varepsilon}}\left[\sup_{\boldsymbol{\theta}\in B[R]}\sum_{i=1}^{n}\varepsilon_{i}\left\langle U\Lambda^{1/2}\boldsymbol{z}_{i},\boldsymbol{\theta}\right\rangle\right]$
$=2R\mathbb{E}_{\boldsymbol{\varepsilon}}\left[\left\|\sum_{i=1}^{n}\varepsilon_{i}\Lambda^{1/2}\boldsymbol{z}_{i}\right\|\right]$
$\leq 2R\mathbb{E}_{\boldsymbol{\varepsilon}}\left[\left\|\sum_{i=1}^{n}\varepsilon_{i}\Lambda^{1/2}\boldsymbol{z}_{i}\right\|^{2}\right]^{1/2}$
$=2R\left(\sum_{i=1}^{n}\left\langle\Lambda\boldsymbol{z}_{i},\boldsymbol{z}_{i}\right\rangle\right)^{1/2}$
and thus
$\mathbb{E}_{\boldsymbol{Z}}\left[\sup_{\boldsymbol{\theta}\in B[R]}G_{t}^{\boldsymbol{\theta}}\left(\boldsymbol{Z}\right)\right]\leq\frac{2R}{n}\sqrt{\sum_{i=1}^{n}\mathbb{E}_{\boldsymbol{Z}}\left[\left\langle\Lambda\boldsymbol{Z}_{i},\boldsymbol{Z}_{i}\right\rangle\right]}\leq 2R\sqrt{\frac{\mathrm{tr}(\Sigma)}{n}}$ (60)
by Jensen’s inequality. Therefore, the statement holds true. ∎